1. Consider the following scenario:

At a price of $20 for a large bag of coffee beans, Trader Sam sells 400 bags. When he raises the price to $22 per bag, he only sells 361 bags. Calculate the elasticity of demand using either the simple method or the midpoint method. IMPORTANT NOTE: Round all percentages to whole numbers in your calculations.

Show your working here (4 points):

The price elasticity of demand is:   

  

This number tells us that the price elasticity of demand for this product is:   

RELATIVELY INELASTIC / EXTREMELY ELASTIC / UNITARY ELASTIC / CLOSE TO UNITARY ELASTIC

How are Trader Sam’s revenues affected by this price change?

HE EARNS A LOT MORE / HE EARNS A LOT LESS / HE EARNS ROUGHLY THE SAME REVENUES

Does Trader Sam benefit from raising the price? YES / NO

Does Trader Sam benefit from lowering the price? YES / NO

How would you expect the demand curve to look?

RELATIVELY STEEP / RELATIVELY FLAT / A DOWNWARD SLOPING CURVE

2. Consider the following scenario:

Trader Sam sells 200 pounds of organic kale at a price of $5 per pound. When he lowers the price to $4.75 per pound, he increases his sales to 250 pounds. Calculate the elasticity of demand using either the simple method or the midpoint method. IMPORTANT NOTE: Round all percentages to whole numbers in your calculations.

Show your working here (4 points):

The price elasticity of demand is:   

  

This number tells us that the price elasticity of demand is:   

INELASTIC / ELASTIC / UNITARY ELASTIC

How are Trader Sam’s revenues affected by this price change?

HE EARNS MORE / HE EARNS LESS / HE EARNS EXACTLY THE SAME REVENUES

Does Trader Sam benefit from lowering the price? YES / NO

Should Trader Sam lower the price? YES / NO

How would you expect the demand curve to look?

RELATIVELY STEEP / RELATIVELY FLAT / A DOWNWARD SLOPING CURVE

3. Consider the following scenario:

Trader Sam sells 100 packs of strawberries at a price of $2.50. When he lowers the price to $2.00 per pound, he increases his sales to 110 packs. Calculate the elasticity of demand using either the simple method or the midpoint method. IMPORTANT NOTE: Round all percentages to whole numbers in your calculations.

Show your working here (4 points):

The price elasticity of demand is:   

  

This number tells us that the price elasticity of demand is:   

INELASTIC / ELASTIC / UNITARY ELASTIC

How are Trader Sam’s revenues affected by this price change?

HE EARNS MORE / HE EARNS LESS / HE EARNS EXACTLY THE SAME REVENUES

Does Trader Sam benefit from lowering the price? YES / NO

Should Trader Sam lower the price? YES / NO

How would you expect the demand curve to look?

RELATIVELY STEEP / RELATIVELY FLAT / A DOWNWARD SLOPING CURVE

BASIC FORMULA: Percentage change in quatity demanded Percentage change in price Price elasticity of demand- SIMPLE METHOD: I (New Quantity - Original Quantity) / Original Quantity ] X 100 Price Elasticity of Demand (New Price - Original Price) / Original Price ] X 100 MIDPOINT METHOD: ((02-01)/I(0+01/211 X100 Price Elasticity of Demand- (P2 - PI) I (P2 + P1)/21 X100
Answer #1

Price elasticity of demand by simple method

Price elasticity(ep)= Percentage change in quantity demanded/Percentage change in price.

Percentage change in price= 22-20/20X100= 2/20X100= 10%

Percentage change in quantity demanded= 400-361/400X100= 39/400X100= 9.75%= 10%(whole number as per instructions given in question)

Price elasticity of demand(ep)= 10/10= 1, ep=1, unitary elasticity of demand.

Price elasticity of demand for this product is unitary means as price increases from 20 to 22 demand for that product also decreases in the same proportion.

Effect on Sam's revenue.

At initial price of $20 and initial quantity demanded 400, he earns $8000 in terms of revenue

At increased price of $22 and decreased quantity demanded 361, he earns $7942 in terms of revenue, he although earns less but since we have taken percentage change in quantity demanded as 10%(whole number) instead of 9.75% so we will assume that he earns exactly the same revenue.

When price falls, suppose price falls from $22 to $20 then quantity demanded increases from 361 to 400 in that case trader SAM does not get benefitted since price elasticity of demand is 0.11 which is less than 1, please refer to below.

Percentage change in Price after fall in price= 22-20/22X100= 2/22X100= 9.09% or 9%(whole number)

Percentage change in quantity demanded= 400-361/361x100= 39/361x100=  1.08% or 1%(whole number)

Price elasticity of demand(ep)= 1/9= 0.11, ep<1

Demand curve would be a downward sloping curve since with a change in price from $20 to $22 of a product its quantity demanded decreases from 400 to 361 in the same proportion.